Evening shift 12 Jan. 2000


Subject: Evening shift 12 Jan. 2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:21:50 CST


                     Shift summary for day shift 12/Jan/2000
On Shift: RW and NS

Hours of beam available at all intensities: 6.75
Hours of beam available at nominal intensity: 6.25
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 5.75

Accounting of lost time: Accelertor problems, power glitches on Kirk road.
========================================================================
Tapes brought to FCC this shift: none
========================================================================
Accesses during shift: none
========================================================================
Runs ended worth looking at this shift: 15493, 15494, 15497
=======================================================================
Overview of shift: Tryed to keep the DA working when there was beam.
========================================================================
Description of Shift:
run spill time comment
------------------------------------------------------------------------
15493 1 15:30 New run started after complete restart of
                        all computers, but no beam from accelerator.
          11 15:41 First spill with beam, low intensity.
          12 Low intensity beam, MI extraction problems,
          77 Still low intensity beam 1.5e11.
          83 17:20 first spill with 2.1e12.
          94 17:36 MCR called, said they have to shut off for
                        next half hour to reboot their computers.
15494 1 17:47 New run started while they were reinitializing
                        the MI computers, beam at 1.5e12.
                17:55 Bob T in consoltation with Ron Ray has lowered
                        the temperature of the block house by 3C out
                        of concerns that the increase in temperature
                        was causing our E/p discrepency. (it was hovering
                        at 21 which is about 4C to high, Bob T. says if
                        we don't get this down then we may have a big
                        problem with the cooling)
          16 Beam went away.
          19 Beam back.
          27 Beam gone.
          29 Beam returns.
          31 Kfrend and reset reg adc.
          46 No beam, Electric glich has knocked out Accelertor
                        RF they have no estimate for recovery, all of our
                        computers are still running and we have to wait
                        for beam to return.
                 18:52 Run stopped.
15497 1 19:30 Run started and beam returns, triggers and DA
                        hangs, RK and Sasha take all the triggers out and
                        then put them back in one spill at a time.
           8 first spill with triggers 1 & 2 back in.
          12 Take MA crate out of the readout.
          15 No beam this spill.
          17 Spill back, off target.
          19 Beam back on target after calling MCR spill
                        structure still kind of bad but better than
                        before.
        20-40 Study of trigger 1 and H-dibaryon on and off
                        to see that the H-diabaryon filter code was the
                        cause of some L3 overload on KTeV2 processors.
                        H-diabaryon trigger was turned off until this
                        problem is understood and solved. Bob T paged
                        leave off and then Arun was paged
                        and is coming in to look into problem.
          47 20:30 Temperature in block house seems to be a problem
                        still, Ed left and if problem persists we are
                        to page Ron Ray after 21:01 hrs.
          74 21:05 run stopped, to try Arun's changes to L3
                        H-dibaryon code.
15498 21:10 New run started, had to reset all the display
                        windows so you have to do a drt-fresh start
                        after this run, this run has new H-dibaryon L3
                        code.
                 21:25 Ron Ray calls back concerning CsI block house high
                        tempeature, suggests we call Leonard.
                 22:55 Ron Ray after working for 30 min and cycling
                        the power on the Bry-air unit he claims the
                        cooling for the blockhouse is not functioning.
                 23:20 Beam quality so bad that traces went off of
                        scale of the storage scope, called MCR and they
                        too agreed that it was horrible.
                 24:00 shift ends, this was not the shortest shift
                        summary ever written!

=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=
Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:21:53 CST